# Finding product sets in some classes of amenable groups

### v1

@@ Abstract.
In [KMRR23b], using methods from ergodic theory, a longstanding conjecture of Erdős (see [Erd73, Page 305]) about sumsets in large subsets of the natural numbers was resolved. In this paper, we extend this result to several important classes of amenable groups, including all finitely generated virtually nilpotent groups, and all abelian groups $(G,+)$ with the property that the subgroup $2G:=\{g+g:g\in G\}$ has finite index. We prove that in any group $G$ from the above classes, any $A\subset G$ with positive upper Banach density contains a shifted product set of the form $\{tb_{i}b_{j}\colon i<j\}$, for some infinite sequence $(b_{n})_{n\in\mathbb{N}}$ and some $t\in G$. In fact, we show this result for all amenable groups that posses a property which we call square absolute continuity. Our results provide answers to several questions and conjectures posed in [KMRR23a].
@@ 1. Introduction
In [KMRR23b], Bryna Kra, Joel Moreira, Florian K. Richter and Donald Robertson, using methods from ergodic theory, proved that every subset $A$ of the positive integers with positive upper Banach density contains $\{b_{1}+b_{2}+t:b_{1}\neq b_{2}\in B\}$ for some infinite set $B\subset A$ and some $t\in\mathbb{N}$. This resolved a longstanding conjecture of Erdős (see [Erd73, Page 305]).
A natural question to ask is whether this result generalizes to other countable groups, such as $\mathbb{Z}^{d}$ for $d\geq 2$ or the discrete Heisenberg group, for example. The purpose of this paper is to extend the result in [KMRR23b] to several important classes of amenable groups, including all finitely generated virtually nilpotent groups and all abelian groups $(G,+)$ with the property that the subgroup consisting of the elements $2g:=g+g$, where $g\in G$, has finite index. To this end, we first extend the result to all amenable groups satisfying a property that we call square absolute continuity (see 1.4). Then we show that our result applies to the aforementioned classes of groups, by showing that they are (virtually) square absolutely continuous. Our main results provide partial answers to some questions and conjectures posed in [KMRR23a] regarding product sets in large subsets of amenable groups.
@@
Given a group $(G,\cdot)$, for any $g,h\in G$, Given a sequence $B=(b_{n})_{n\in\mathbb{N}}\subset G$, we define
@@
Let $G$ be an amenable group and $\phi:G\to G$ be a map. We say that $G$ is $\phi$-absolutely continuous if $G$ admits two Følner sequences $\Phi=(\Phi_{N})_{N\in\mathbb{N}}$ and $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ satisfying the following: for any $\varepsilon>0$ there exists some $\delta>0$ such that for any $u:G\to[0,1]$ satisfying
@@ 1.1. Main results
The first main theorem of this paper is the following:
@@
Theorem 1.5 provides a positive answer to [KMRR23a, Question 5.16], with the extra assumption that $G$ is square absolutely continuous, and under the weaker assumption that the set $A$ has positive left upper Banach density, instead of positive upper Banach density.
@@
Note that Theorem 1.5 immediately implies an analogous result for right upper Banach density instead of left upper Banach density. Indeed, through the map $g\mapsto g^{-1}$, Theorem 1.5 is equivalent to the assertion that for any $A\subset G$ with positive right upper Banach density, there exists an infinite sequence $C=(c_{n})_{n\in\mathbb{N}}\subset A$ and some $r\in G$ such that
@@
  • $G$ is nilpotent if its lower central series is finite, that is to say
@@
  • $G$ is torsion-free if it does not have any non-trivial element of finite rank, that is to say, for any $g\in G$ with $g\neq e_{G}$ and any $n\in\mathbb{N}$, we have $g^{n}\neq e_{G}$.
@@
In any finitely generated nilpotent group $G$, there exist some $s\in\mathbb{N}$ (depending on the degree of nilpotency and the number of generators of $G$), some $a_{i}\in G$ and some functions $t_{i}:G\to\mathbb{Z}$, for $1\leq i\leq s$, such that any $x\in G$ can be written as $x=a_{1}^{t_{1}(x)}\cdots a_{s}^{t_{s}(x)}$. The $s$-tuple $(a_{1},\dots,a_{s})$ is a Mal’cev basis and the $s$-tuple $(t_{1},\dots,t_{s})$ is a Mal’cev coordinate system with respect to this Mal’cev basis. If $G$ is also torsion-free, then the coordinate maps are injective and hence we can identify $G$ with $\mathbb{Z}^{s}$ and it is convenient to also identify any $x\in G$ with its coordinates $(t_{1}(x),\dots,t_{s}(x))\in\mathbb{Z}^{s}$. The above facts about Mal’cev bases can be found in [KM79, Chapter 17.2].
@@
Combining Theorems 1.5 and 1.7 we have that every torsion-free finitely generated nilpotent group satisfies the conclusion of Theorem 1.5. In fact, we prove the following slight strengthening:
@@
In particular, this holds for the group $UT(n,F)$111 $UT(n,F)$ is the unitriangular $n\times n$ matrix group with entries from $F$. Note that $U(3,\mathbb{Z})$ is the well-known $3\times 3$ Heisenberg group., where $n\in\mathbb{N}$ and $F$ is any infinite, finitely generated field.
@@
Note that the class of finitely generated virtually nilpotent groups coincides, in view of Gromov’s theorem, with the class of finitely generated groups of polynomial growth.
@@
In [KMRR23a, Conjecture 5.14], it is conjectured that in any countable abelian group $G$, every set of positive upper Banach density contains a set of the form $B\oplus B+t=\{b_{1}+b_{2}+t:b_{1}\neq b_{2}\in B\}$ for an infinite set $B\subset G$ and some $t\in G$. It follows from Corollary 1.9 that this conjecture holds under the additional assumption that $G$ is finitely generated, and moreover, it extends [KMRR23b, Theorem 1.2] from $(\mathbb{N},+)$ to all finitely generated abelian groups. In fact, we are able to verify [KMRR23a, Conjecture 5.14], and hence extend [KMRR23b, Theorem 1.2] to an even larger collection of abelian groups, that contains all the finitely generated abelian groups along with some infinitely generated ones. The following theorem allows us to do so:
@@
The following corollary is an obvious consequence of Theorems 1.5 and 1.11:
@@
In the semigroup $(\mathbb{N},\times)$ the sub-semigroup $\{n^{2}\colon n\in\mathbb{N}\}$ has infinite index (and zero density), and unfortunately, our methods do not allow us to handle this case.
@@ 1.2. More product sets and open questions
True (Theorem 1.5; for $\overline{d}_{\Phi}(A)>0$ for some left Følner $\Phi$)
@@
True (Remark 1.13; for $\overline{d}_{\Phi}(A)>0$ for some left Følner $\Phi$)
@@
False (Example 6.1; with $\overline{d}_{\Phi}(A)>0$ for some left Følner $\Phi$)
@@
False (Example 6.2; with $\overline{d}_{\Phi}(A)=1$ for some left Følner $\Phi$)
@@
False (Example 6.3; with $\overline{d}_{\Phi}(A)=1$ for some left Følner $\Phi$)
@@
It remains interesting to ask whether product sets of the form $B\odot B$ can be found in sets with positive upper Banach density. Unfortunately, our methods here are insufficient to handle this case. In this spirit, we conclude this section with the two questions below. We remark that the second one is a special case of [KMRR23a, Question 5.17].
Question 1.14.
Let $G$ be a square absolutely continuous group and $A\subset G$ be a set of positive upper Banach density. Is it true that there exists some set $B\subset G$ such that
@@
Question 1.15.
Let $G$ and $A$ be as in 1.14. Is it true that there exists some set $B\subset G$ such that
@@ 1.3. Proof ideas
To prove Theorem 1.5, we follow an ergodic-theoretic approach and we employ ideas similar to the ones used in [KMRR23b] in the setting of $(\mathbb{N},+)$. This approach is based on methods that were introduced in [KMRR22] to generalize another sumset conjecture of Erdős, which was initially proved in [MRR19] by Moreira, Richter and Robertson. However, the generality of the setting of amenable groups compared to $(\mathbb{N},+)$ causes several issues and complications that we need to handle differently. These issues, along with the new ideas we develop to deal with them, are briefly discussed below.
@@
One of the main obstructions we had to overcome in our proof is the lack of commutativity of $G$. The most notable among the issues that this leads to is that the Kronecker factor does not have the structure of an abelian group, but instead, it is a homogeneous space $Z=K/H$, for some compact group $K$. This makes the study of Erdős progressions more technically challenging. To be more precise, the abelian nature of the Kronecker factor in the setting of $(\mathbb{N},+)$-actions is heavily used in [KMRR23b]. Consequently, due to the absence of commutativity in our case, many of the techniques in [KMRR23b] do not generalize easily to our setting. Another difficulty that arises in non-commutative groups is the erratic behavior of the set of squares $G^{2}$. In particular, orbits of points along $G^{2}$ may be trapped in zero-measure regions, which causes serious trouble in finding Erdős progressions. The assumption that $G$ is square absolutely continuous is critical in avoiding this scenario. In addition, we need an extension of a result of Host and Kra ([HK09, Proposition 6.1]) concerning actions of $(\mathbb{N},+)$, to the more general setting of amenable group actions (see Lemma 3.5, proof in Appendix A).
Acknowledgements. We would like to thank Joel Moreira and Florian K. Richter for their insightful suggestions, beneficial comments and constant support throughout the writing of this paper. We also want to thank Ethan Ackelsberg and Felipe Hernández Castro for their useful comments on earlier drafts of this paper.
The first author gratefully acknowledges support from the Swiss National Science Foundation grant TMSGI2-211214. The second author is supported by the Warwick Mathematics Institute Centre for Doctoral Training, and gratefully acknowledges funding by University of Warwick’s Chancellors’ International Scholarship scheme.
@@ 2. Preliminaries
Fix a topological $G$-system $(X,T)$. Let $M(X)$ denote the space of Borel probability measures on $X$, equipped with the ${}^{\ast}$ topology, which is compact and metrizable. A measure $\mu\in M(X)$ is said to be $T$-invariant, if it is invariant under $T_{g}$ for all $g\in G$. The subset of $M(X)$ consisting of $T$-invariant measures is denoted by $M^{T}(X)$, and it is a closed and convex subset of $M(X)$. The Borel $\sigma$-algebra on $X$ is denoted by $\mathscr{B}_{X}$ or just $\mathscr{B}$, if no confusion may arise.
For $\mu\in M^{T}(X)$, the action $T$ on the Borel probability space $(X,\mu)$ is called a measure-preserving $G$-action and $(X,\mu,T)$ is called a measure-preserving $G$-system. Note that we omit writing the symbol for the $\sigma$-algebra, and from now on, whenever this happens, the implied $\sigma$-algebra will be the Borel. For simplicity, we refer to the above as $G$-actions, and $G$-systems, respectively.
@@
We remark that we are only considering $G$-systems where $G$ acts on the prescribed space from the left, and then any associated Følner will be considered left, without mentioning it, unless it is necessary.
@@
Theorem 2.1 (Mean Ergodic Theorem for $G$-systems, see [Gla03, Theorem 3.33]).
@@
Theorem 2.2 (Disintegration of measures, see [HK18, Chapter 2, Section 2.5]).
@@
  • $\mu_{x}([x]_{\mathscr{D}})=1$, where $[x]_{\mathscr{D}}=\cap_{x\in D\in\mathscr{D}}D.$
@@
In this paper, we will extensively make use of disintegrations over (continuous) factor maps. Let $\pi:(X,\mu,T)\to(Y,\nu,S)$ be a factor map between two $G$-systems. In the setting of Theorem 2.2, let $\mathscr{D}=\pi^{-1}(\mathscr{B}_{Y})$, where $\mathscr{B}_{Y}$ is the Borel $\sigma$-algebra on $Y$. Then Theorem 2.2 gives a disintegration $y\mapsto\mu_{y}$ defined on $Y$, which is unique up to $\nu$-null measure sets, and satisfies the following:
  • for $\nu$-almost every $y\in Y$,
    (2.1) $\mathbb{E}_{\mu}(f\>|\>Y)(y):=\mathbb{E}_{\mu}(f\>|\>\mathscr{D})(y)=\int_{Y}f~{}\mathrm{d}\mu_{y},$
@@
Theorem 2.3 (Ergodic decomposition of $G$-systems, see [Gla03, Theorem 3.22]).
@@
where $\delta_{x}$ is the Dirac mass at $x\in X$ and the limit is in the ${}^{\ast}$ topology. If $a$ is generic for $\mu$ along $\Phi$, then we denote this by $a\in\texttt{{gen}}(\mu,\Phi)$.
@@
Proposition 2.6 (see [Lin99, Theorem 1.2 and Proposition 1.4]).
Let $(X,\mu,T)$ be a system and $\Phi$ be a Følner sequence in $G$. Then there is a subsequence $\Psi$ of $\Phi$ such that for all $f\in L^{1}(\mu)$,
@@
The next two lemmas follow easily from Proposition 2.6:
@@
Finally, it will be useful to have the following generalisation of [Fur81, Proposition 3.9], whose proof is again the same as for actions of $(\mathbb{N},+)$, but we include it for the convenience of the reader.
@@
Since $x_{0}\in\overline{\mathcal{O}_{T}(a)}$, there exists some $(g_{n})_{n\in\mathbb{N}}\subset G$ such that $T_{g_{n}}a\to x_{0}$, so that we may assume that the equation above holds if we substitute $x_{0}$ with $T_{g_{n}}a$. Consider the Følner sequence $\Psi=(\Psi_{n})$ given by $\Psi_{n}=\Phi_{N_{n}}g_{n}$. It follows that for every $n\in\mathbb{N}$ and any $j=1,2,\dots,n$,
@@
Theorem 2.10 (Jacobs-de Leeuw-Glicksberg decomposition, see [KL16, Theorem 2.24]).
@@
Now we will give a description of the factor of $(X,\mu,T)$ corresponding to the subspace $\mathcal{H}_{\textup{c}}$ of $L^{2}(X)$. Note that if $f\in\mathcal{H}_{\textup{c}}$, then for all $g\in G$, $T_{g}f\in\mathcal{H}_{\textup{c}}$, so $\mathcal{H}_{\textup{c}}$ in invariant under the action of $T$ on $L^{2}(X)$. Let $\mathcal{A}$ be the smallest $\sigma$-algebra with respect to which all functions in $\mathcal{H}_{\textup{c}}$ are measurable. Then $\mathcal{A}$ is $T$-invariant $\sigma$-algebra contained in the Borel $\sigma$-algebra of $X$. Therefore, the system $(X,\mathcal{A},\mu,T)$ is a factor of the original system, with the factor map being the identity $\textup{id}:X\to X$. This factor is called the Kronecker factor of $(X,\mu,T)$.
Our goal now is to give a nice algebraic description of the Kronecker factor when the $G$-system $(X,\mu,T)$ is ergodic.

### v3

@@ Abstract.
In [15], using methods from ergodic theory, a longstanding conjecture of Erdős (see [5, Page 305]) about sumsets in large subsets of the natural numbers was resolved. In this paper, we extend this result to several important classes of amenable groups, including all finitely generated virtually nilpotent groups, and all abelian groups $(G,+)$ with the property that the subgroup $2G:=\{g+g:g\in G\}$ has finite index. We prove that in any group $G$ from the above classes, any $A\subset G$ with positive upper Banach density contains a shifted product set of the form $\{tb_{i}b_{j}\colon i<j\}$, for some infinite sequence $(b_{n})_{n\in\mathbb{N}}$ and some $t\in G$. In fact, we show this result for all amenable groups that posses a property which we call square absolute continuity. Our results provide answers to several questions and conjectures posed in [13].
1991 Mathematics Subject Classification:
Primary: 05D10, 37A15; Secondary: 11B13, 11B30.
@@ 1. Introduction
In [15], Bryna Kra, Joel Moreira, Florian K. Richter and Donald Robertson, using methods from ergodic theory, proved that every subset $A$ of the positive integers with positive upper Banach density contains $\{b_{1}+b_{2}+t:b_{1}\neq b_{2}\in B\}$ for some infinite set $B\subset A$ and some $t\in\mathbb{N}$. This resolved a longstanding conjecture of Erdős (see [5, Page 305]).
A natural question to ask is whether this result generalizes to other countable groups, such as $\mathbb{Z}^{d}$ for $d\geq 2$ or the discrete Heisenberg group, for example. The purpose of this paper is to extend the result in [15] to several important classes of amenable groups, including all finitely generated virtually nilpotent groups and all abelian groups $(G,+)$ with the property that the subgroup consisting of the elements $2g:=g+g$, where $g\in G$, has finite index. To this end, we first extend the result to all amenable groups satisfying a property that we call square absolute continuity (see 1.4). Then we show that our result applies to the aforementioned classes of groups, by showing that they are (virtually) square absolutely continuous. Our main results provide partial answers to some questions and conjectures posed in [13] regarding product sets in large subsets of amenable groups.
@@
Given a group $(G,\cdot)$ and a sequence $B=(b_{n})_{n\in\mathbb{N}}\subset G$, we define
@@
Let $G$ be an amenable group and $\phi:G\to G$ be a map. We say that $G$ is $\phi$-absolutely continuous if $G$ admits two left Følner sequences $\Phi=(\Phi_{N})_{N\in\mathbb{N}}$ and $\Psi=(\Psi_{N})_{N\in\mathbb{N}}$ satisfying the following: for any $\varepsilon>0$ there exists some $\delta>0$ such that for any $u:G\to[0,1]$ satisfying
@@
Remark 1.5.
For a set $C\subset G$, we denote by $C^{-1}$ the set $C^{-1}:=\{g^{-1}:g\in C\}$. Note that when $\phi$ is the squaring map $\phi=s_{G}$, the existence of two left Følner sequences in 1.4 is equivalent to the existence of two right Følner sequences: for any pair $\Phi$, $\Psi$ of left Følner sequences, the pair $\Phi^{-1}=(\Phi_{N}^{-1})_{N\in\mathbb{N}}$, $\Psi^{-1}=(\Psi_{N}^{-1})_{N\in\mathbb{N}}$ is a pair of right Følner sequences, and if $\Phi$, $\Psi$ satisfy the conditions of 1.4, then so do $\Phi^{-1}$, $\Psi^{-1}$. More precisely, given $\varepsilon>0$, there is some $\delta>0$ so that the conditions of 1.4 are satisfied for $\Phi,\Psi$. Let $u:G\to[0,1]$ satisfy $\limsup_{N\to\infty}\frac{1}{|\Phi_{N}^{-1}|}\sum_{g\in\Phi_{N}^{-1}}u(g)<\delta$ and consider the map $w:G\to[0,1],\>w(g):=u(g^{-1})$. Then $\limsup_{N\to\infty}\frac{1}{|\Phi_{N}|}\sum_{g\in\Phi_{N}}w(g)<\delta$, and therefore $\limsup_{N\to\infty}\frac{1}{|\Psi_{N}|}\sum_{g\in\Psi_{N}}w(g^{2})<\varepsilon$, which in turn implies that $\limsup_{N\to\infty}\frac{1}{|\Psi_{N}^{-1}|}\sum_{g\in\Psi_{N}^{-1}}u(g^{2})<\varepsilon$.
@@ 1.1. Main results
Throughout, we say that a sequence $(b_{n})_{n\in\mathbb{N}}$ in $G$ is infinite if the set $\{b_{n}\colon n\in\mathbb{N}\}$ is infinite. The first main theorem of this paper is the following:
@@
Theorem 1.6 provides a positive answer to [13, Question 5.16], with the extra assumption that $G$ is square absolutely continuous, and under the weaker assumption that the set $A$ has positive left upper Banach density, instead of positive upper Banach density.
@@
Note that Theorem 1.6 immediately implies an analogous result for right upper Banach density instead of left upper Banach density. Indeed, since, by Remark 1.5, square absolute continuity is preserved through the map $g\mapsto g^{-1}$, Theorem 1.6 is equivalent to the assertion that for any $A\subset G$ with positive right upper Banach density, there exists an infinite sequence $C=(c_{n})_{n\in\mathbb{N}}\subset A$ and some $r\in G$ such that
@@
  • $G$ is nilpotent if its lower central series is finite, that is to say there is $n\in\mathbb{N}$ such that
@@
  • $G$ is torsion-free if it does not have any non-trivial element of finite order, that is to say, for any $g\in G$ with $g\neq e_{G}$ and any $n\in\mathbb{N}$, we have $g^{n}\neq e_{G}$.
@@
In any finitely generated nilpotent group $G$, there exist some $s\in\mathbb{N}$ (depending on the degree of nilpotency and the number of generators of $G$), some $a_{i}\in G$ and some functions $t_{i}:G\to\mathbb{Z}$, for $1\leq i\leq s$, such that any $x\in G$ can be written as $x=a_{1}^{t_{1}(x)}\cdots a_{s}^{t_{s}(x)}$. The $s$-tuple $(a_{1},\dots,a_{s})$ is a Mal’cev basis and the $s$-tuple $(t_{1},\dots,t_{s})$ is a Mal’cev coordinate system with respect to this Mal’cev basis. If $G$ is also torsion-free, then the coordinate maps are injective and hence we can identify $G$ with $\mathbb{Z}^{s}$ and it is convenient to also identify any $x\in G$ with its coordinates $(t_{1}(x),\dots,t_{s}(x))\in\mathbb{Z}^{s}$. The above facts about Mal’cev bases can be found in [11, Chapter 17.2].
@@
Combining Theorems 1.6 and 1.8 we have that every torsion-free finitely generated nilpotent group satisfies the conclusion of Theorem 1.6. In fact, we prove the following slight strengthening:
@@
In particular, this holds for the group $UT(n,F)$111 $UT(n,F)$ is the unitriangular $n\times n$ matrix group with entries from $F$. Note that $U(3,\mathbb{Z})$ is the well-known $3\times 3$ Heisenberg group., where $n\in\mathbb{N}$ and $F=(F,+,\cdot)$ is any infinite commutative unital ring with the property that the additive group $(F,+)$ is finitely generated.
@@
Note that the class of finitely generated virtually nilpotent groups coincides, in view of Gromov’s theorem [8], with the class of finitely generated groups of polynomial growth.
@@
In [13, Conjecture 5.14], it is conjectured that in any countable abelian group $G$, every set of positive upper Banach density contains a set of the form $B\oplus B+t=\{b_{1}+b_{2}+t:b_{1}\neq b_{2}\in B\}$ for an infinite set $B\subset G$ and some $t\in G$. It follows from Corollary 1.10 that this conjecture holds under the additional assumption that $G$ is finitely generated, and moreover, it extends [15, Theorem 1.2] from $(\mathbb{N},+)$ to all finitely generated abelian groups. In fact, we extend [15, Theorem 1.2] to an even larger collection of abelian groups, that contains all the finitely generated abelian groups along with some infinitely generated ones. The following theorem allows us to do so:
@@
The following corollary is an obvious consequence of Theorems 1.6 and 1.12:
@@
Corollary 1.13 is in fact optimal, in the sense that $2G$ being a finite-index subgroup of $G$ is a necessary assumption. As shown in a recent paper of Ethan Ackelsberg [1], if $2G$ has infinite index in $G$, then one can always find a set $A$ with upper Banach density arbitrarily close to $1$ which does not contain any shifted sumset $t+B\oplus B$ of some infinite set $B$. Therefore, Corollary 1.13 along with the work in [1] fully resolve [13, Conjecture 5.14].
@@ 1.2. More product sets and open questions
It remains interesting to ask whether product sets of the form $B\odot B$ can be found in sets with positive upper Banach density. Unfortunately, our methods here are insufficient to handle this case. In this spirit, we conclude this section with the two questions below. We remark that the second one is a special case of [13, Question 5.17].
Question 1.15.
Let $G$ be a square absolutely continuous group and $A\subset G$ be a set of positive upper Banach density. Is it true that there exists some infinite set $B\subset G$ such that
@@
Question 1.16.
Let $G$ and $A$ be as in 1.15. Is it true that there exists some infinite set $B\subset G$ such that
@@ 1.3. Proof ideas
To prove Theorem 1.6, we follow an ergodic-theoretic approach and we employ ideas similar to the ones used in [15] in the setting of $(\mathbb{N},+)$. This approach is based on methods that were introduced in [14] to generalize another sumset conjecture of Erdős, which was initially proved in [18] by Moreira, Richter and Robertson. However, the generality of the setting of amenable groups compared to $(\mathbb{N},+)$ causes several issues and complications that we need to handle differently. These issues, along with the new ideas we develop to deal with them, are briefly discussed below.
@@
One of the main obstructions we had to overcome in our proof is the lack of commutativity of $G$. The most notable among the issues that this leads to is that the Kronecker factor does not have the structure of an abelian group, but instead, it is a homogeneous space $Z=K/H$, for some compact group $K$. This makes the study of Erdős progressions more technically challenging. To be more precise, the abelian nature of the Kronecker factor in the setting of $(\mathbb{N},+)$-actions is heavily used in [15]. Consequently, due to the absence of commutativity in our case, many of the techniques in [15] do not generalize easily to our setting. Another difficulty that arises in non-commutative groups is the erratic behavior of the set of squares $G^{2}$. In particular, orbits of points along $G^{2}$ may be trapped in zero-measure regions, which causes serious trouble in finding Erdős progressions. The assumption that $G$ is square absolutely continuous is critical in avoiding this scenario. In addition, we need an extension of a result of Host and Kra ([9, Proposition 6.1]) concerning actions of $(\mathbb{N},+)$, to the more general setting of amenable group actions (see Lemma 3.5, proof in Appendix A).
Acknowledgements. We would like to thank Joel Moreira and Florian K. Richter for their insightful suggestions, beneficial comments and constant support throughout the writing of this paper. We also want to thank Ethan Ackelsberg, Felipe Hernández Castro and the anonymous referee for their useful comments.
The first author gratefully acknowledges support from the Swiss National Science Foundation grant TMSGI2-211214. The second author was supported by the Warwick Mathematics Institute Centre for Doctoral Training, and gratefully acknowledges funding by University of Warwick’s Chancellors’ International Scholarship scheme.
@@ 2. Preliminaries
Fix a topological $G$-system $(X,T)$. Let $M(X)$ denote the space of Borel probability measures on $X$, equipped with the weak topology, which is compact and metrizable. A measure $\mu\in M(X)$ is said to be $T$-invariant, if it is invariant under $T_{g}$ for all $g\in G$. Amenability of $G$ implies that there are $T$-invariant measures in $M(X)$. The subset of $M(X)$ consisting of $T$-invariant measures is denoted by $M^{T}(X)$, and it is a non-empty closed and convex subset of $M(X)$. The Borel $\sigma$-algebra on $X$ is denoted by $\mathscr{B}_{X}$ or just $\mathscr{B}$, if no confusion may arise.
For $\mu\in M^{T}(X)$, the action $T$ on the Borel probability space $(X,\mu)$ is called a measure-preserving $G$-action and $(X,\mu,T)$ is called a measure-preserving $G$-system. Note that we omit writing the symbol for the $\sigma$-algebra, and from now on, whenever this happens, the implied $\sigma$-algebra will be the Borel $\sigma$-algebra. For simplicity, we refer to the above as $G$-actions, and $G$-systems, respectively. Recall that all $G$-actions considered throughout are continuous.
@@
We remark that we are only considering $G$-systems where $G$ acts on the prescribed space from the left, and then any associated Følner sequence will be considered left, without mentioning it, unless it is necessary.
Note that we could define $G$-systems more generally as follows: a $G$-system is a quadruple $(X,\mathcal{A},\mu,T)$, where $X$ is any set, $\mathcal{A}$ is a $\sigma$-algebra on $X$, $\mu$ is a probability measure on $(X,\mathcal{A})$ and $T$ is a left action of $G$ on $X$ which is measurable and preserves $\mu$. We chose to not define systems in that generality, as for our purposes we will always work with the more specific $G$-systems defined above. The only occasion where we need this more general definition of $G$-systems is when we define the Kronecker factor right after Theorem 2.10, in which case $\mathcal{A}$ is a sub-$\sigma$-algebra of the Borel $\sigma$-algebra.
@@
Theorem 2.1 (Mean Ergodic Theorem for $G$-systems, see [7, Theorem 3.33]).
@@
Theorem 2.2 (Disintegration of measures, see [10, Chapter 2, Section 2.5]).
@@
  • For $(\mathscr{D},\mu)$-almost every $x\in X$, $\mu_{x}([x]_{\mathscr{D}})=1$, where $[x]_{\mathscr{D}}=\cap_{x\in D\in\mathscr{D}}D.$
@@
In this paper, we will extensively make use of disintegrations over (continuous) factor maps. Let $\pi:(X,\mu,T)\to(Y,\nu,S)$ be a factor map between two $G$-systems. Given a function $f\in L^{1}(X,\mu)$, the conditional expectation $\mathbb{E}_{\mu}(f\>|\>Y)$ of $f$ with respect to the factor $Y$ is the function in $L^{1}(Y,\nu)$ defined by
$\int_{B}\mathbb{E}_{\mu}(f\>|\>Y)(y)~{}\mathrm{d}\nu(y)=\int_{\pi^{-1}(B)}f(x)~{}\mathrm{d}\mu(x)$
for every Borel measurable set $B\subset Y$. Then Theorem 2.2 gives a disintegration $y\mapsto\mu_{y}$ defined on $Y$, which is unique up to $\nu$-null measure sets, and satisfies the following:
  • for every $f\in L^{1}(X,\mu)$, for $\nu$-almost every $y\in Y$,
    (2.1) $\mathbb{E}_{\mu}(f\>|\>Y)(y)=\int_{X}f~{}\mathrm{d}\mu_{y},$
@@
Theorem 2.3 (Ergodic decomposition of $G$-systems, see [7, Theorem 3.22]).
@@
where $\delta_{x}$ is the Dirac mass at $x\in X$ and the limit is in the weak topology. If $a$ is generic for $\mu$ along $\Phi$, then we denote this by $a\in\texttt{{gen}}(\mu,\Phi)$.
@@
Proposition 2.6 (see [16, Theorem 1.2 and Proposition 1.4]).
Let $(X,\mu,T)$ be a $G$-system and $\Phi$ be a Følner sequence in $G$. Then there is a subsequence $\Psi$ of $\Phi$ such that for all $f\in L^{1}(\mu)$,
@@
The next two lemmas follow easily from Proposition 2.6 and (2.2).
@@
Finally, it will be useful to have the following generalization of [6, Proposition 3.9], whose proof is again the same as for actions of $(\mathbb{N},+)$, but we include it for the convenience of the reader.
@@
Since $x_{0}\in\overline{\mathcal{O}_{T}(a)}$, there exists some $(g_{n})_{n\in\mathbb{N}}\subset G$ such that $T_{g_{n}}a\to x_{0}$, so that we may assume that the equation above holds if we substitute $x_{0}$ with $T_{g_{n}}a$. Consider the Følner sequence $\Psi=(\Psi_{n})$ given by $\Psi_{n}=\Phi_{N_{n}}g_{n}$. Note that this is still a left Følner sequence. It follows that for every $n\in\mathbb{N}$ and any $j=1,2,\dots,n$,
@@
Theorem 2.10 (Jacobs-de Leeuw-Glicksberg decomposition, see [12, Theorem 2.24]).
@@
Now we will give a description of the factor of $(X,\mu,T)$ corresponding to the subspace $\mathcal{H}_{\textup{c}}$ of $L^{2}(X)$. Note that if $f\in\mathcal{H}_{\textup{c}}$, then for all $g\in G$, $T_{g}f\in\mathcal{H}_{\textup{c}}$, so $\mathcal{H}_{\textup{c}}$ in invariant under the action of $T$ on $L^{2}(X)$. Let $\mathcal{A}$ be the smallest $\sigma$-algebra with respect to which all functions in $\mathcal{H}_{\textup{c}}$ are measurable. Then $\mathcal{A}$ is a $T$-invariant $\sigma$-algebra contained in the Borel $\sigma$-algebra of $X$. Therefore, the system $(X,\mathcal{A},\mu,T)$ is a factor of the original system, with the factor map being the identity $\textup{id}:X\to X$. This factor is called the Kronecker factor of $(X,\mu,T)$.
Our goal now is to give a nice algebraic description of the Kronecker factor when the $G$-system $(X,\mu,T)$ is ergodic, but first we need the following definition.
Definition 2.11.
Let $K$ be a compact group, $H$ be a closed subgroup of $K$, and $\alpha\colon G\to K$ be a group homomorphism. Consider the homogeneous space $Z=K/H$. Let also $m$ be the normalized Haar measure on $Z$, and $R=(R_{g})_{g\in G}$, where for each $g\in G$, $R_{g}\colon Z\to Z$ is given by $R_{g}(z)=\alpha(g)z$. Then the $G$-system $(Z,m,R)$ is called a rotation on the homogeneous space $Z$ by $\alpha$.